the number of hours a boat will take to travel 10km upstream if the rate of the boat in still water is x kph and the current is 2.5 kph

Answer by stanbon(48558)

the number of hours a boat will take to travel 10km upstream if the rate of the boat in still water is x kph and the current is 2.5 kph
------------------------
Upstream DATA:
Rate = (x+2.5)kph ; distance = 10 km
EQUATION:
Time = d/r = 10/(x+2.5) hrs.
=====================================
Cheers,
Stan H.